Adv. Math. Econ. 7, 95-111 (2005)

Real indeterminacy of equilibria with real and nominal assets

Ryo Nagata

School of Political Science and Economics, Waseda University 1-6-1 Nishiwaseda, Shinjuku-ku, Tokyo 169-8050, Japan

Received: July 14, 2004 Revised: September 21, 2004

JEL classification: C 60, D 52

Mathematical Subject Classification (2000): 57N15, 58A05, 91B50, 91B52

Abstract. This paper investigates the real indeterminacy of equilibria in an incomplete market model in which there are two periods, with uncertainty in the second, and both real and nominal assets exist. As is well known, the equilibria of a model with real assets behave very differently from the equilibria of a model with nominal assets. Then, what happens if real and nominal assets coexist? In this paper it is demonstrated within a general framework that regardless of the presence of real assets there is generically still real indeterminacy of equilibria whose degree is the same as without the real assets.

Key words: incomplete market, real asset, nominal asset, real indeterminacy.